OpenStudy (anonymous):

If someone could help me get started with this one please? Find the slope of the line that passes through the pair of points (-a+1, b-1) and (a+1, -b)

OpenStudy (anonymous):

I know to use y=mx+b, but where would I start with this one?

OpenStudy (anonymous):

slope=(y_step)/(x_step) so here slope=(-b-(b-1))/(a+1-(-a+1))=1/2a

OpenStudy (anonymous):

Thanks, give me a few minutes and I'll post what I come up with.

OpenStudy (anonymous):

\[m=\frac{ 2b-1 }{ 2a+1 }\]

OpenStudy (anonymous):

I can show my work if need be...

OpenStudy (anonymous):

the y_step is -b-(b-1)=1 the x_step is a+1-(-a+1)=2a

OpenStudy (anonymous):

So you're saying that the answer is \[\frac{ 1 }{ 2a }\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that isn't one of the answer choices...

OpenStudy (anonymous):

what choices are there?

OpenStudy (anonymous):

give me a little bit, I type slow and there are 5 choices...lol

OpenStudy (anonymous):

ah. I'm totally stupid :D it's (-2b+1)/2a

OpenStudy (anonymous):

\[m=\frac{ 1-2b }{ 2a }, m=\frac{ 2b-1 }{ 2a }, m=\frac{ 2b }{ 1-2a }, m=\frac{ 2a }{ 1-2b }, m=\frac{ 1-2b }{ 2a+1 }\]

OpenStudy (anonymous):

so, A is the correct answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... i went wrong big time somewhere, because i came up with B

OpenStudy (anonymous):

and I just saw where I went wrong...lol

OpenStudy (anonymous):

there are 2 points, A and B. let the coordinates of A be x1 and y1, and the coordinates of B be x2 and y2. the slope is (y2-y1)/(x2-x1) or (y1-y2)/(x1-x2) but (y2-y1)/(x1-x2) is not the same, it's the opposite ;)

OpenStudy (anonymous):

yeah, I forgot basic algebra...lol here is what I initially came up with using y=mx+b: \[m=\frac{ b }{ a+1 }-\frac{ b-1 }{ -a+1 }=\frac{ 2b-1 }{ 2a }\]

OpenStudy (anonymous):

Once I use basic algebra, I came up with (still using y=mx+b) \[m=\frac{ 1-2b }{ 2a }\]

OpenStudy (anonymous):

good for you :)

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

you are welcome